Circular No. 7879

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ch IN ANONYMOUS GALAXY T. Puckett and A. Langoussis, Mountain Town, GA, report the discov-

ery of an apparent supernova (mag 18.4) on an unfiltered CCD frame taken with the Puckett Observatory 0.35-m automated supernova patrol telescope on Apr. 16.14 UT (and confirmed on a frame taken on Apr. 17.08). The new object is located at α = 11^{h}14^{m}31^{s}.72, δ = +33^{o}48′18′′.3 (equinox 2000.0), which is 4′′.5 west and 3′′.0 south of the center of its apparent host galaxy. SN 2002ch is not present on images taken by Puckett on 2001 Feb. 25, Apr. 10, May 9, and Nov. 20 (limiting mag ∼ 20.5) or on Palomar Sky Survey plates taken on 1990 Mar. 19 and 1992 Feb. 4 (limiting mag ∼ 21.0) and on 1953 Mar. 9 (limiting mag ∼ 20.0).

POSSIBLE NOVA IN SAGITTARIUS W. Liller, Viña del Mar, Chile, reports his measurement for the po-

sition of the possible nova reported on IAUC 7878: α = 17^{h}59^{m}59^{s}.38 ± 0^{s}.08, δ = −30^{o}53′21′′.9 ± 1′′.0 (equinox 2000.0). On Apr. 16.40 UT, a lowdispersion objective-prism CCD spectrogram, taken by Liller with a 0.20-m Schmidt camera, showed the object to have narrow Hα in emission with a strength ∼ 2.5 times that of the surrounding continuum (FWHM ≈ 800 km/s); Hβ is present but very faint, suggesting that the object is considerably reddened. The [Fe II] multiplet-42 lines also seem to be present. On Apr. 16.3814, Liller measured the star’s magnitude to be V = 9.84. T. Kato, Kyoto University, reports that K. Haseda (Aichi, Japan) found nothing to magnitude 13.1 on a T-Max 400 film taken on Apr. 12.79.

V4334 SAGITTARII F. Kerber, N. Pirzkal, and M. R. Rosa, Space Telescope European

Coordinating Facility, European Southern Observatory (ESO), Garching, report the discovery of emission lines from V4334 Sgr, indicating that the star has again left the asymptotic-giant branch (AGB) only a few years after its final helium flash in 1996. Spectra obtained with ESO’s Very Large Telescope (+ FORS2) on 2001 June 17 and 22 show [N II] 654.8- and 658.3-nm, as well as [O II] 732.5-nm and [O I] 630.0-nm. Hα is marginally detected at ≤ 20 percent of [N II] 658.3-nm. Kerber et al. conclude that the star has started to ionize the H-deficient matter in its surrounding dusty envelope of ejecta, indicating that the star’s temperature has increased from ∼ 5500 K in 1997 to ≥ 20 000 K, and that the star is now retracing its post-AGB evolution toward the planetary-nebula phase for a second time.

2002 April 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green